OpenStudy (anonymous):

Graph each system of inequalities 5x+y >2 2x +y < 1 -

OpenStudy (anonymous):

These are tedious. But, can be done. Graph the lines first as if they are "equalities". Then, shade the region below or above that line depending on how the inequality works. All points in the shaded region satisfy.

OpenStudy (anonymous):

can you help me with it.. please

OpenStudy (anonymous):

Let us take 2x + y < 1. Treat it for now as 2x + y = 1. x intercept is when y = 0. So, 2x = 1 => x = 1/2 y intercept is when x = 0. So, y = 1. Now, draw this line as follows using the x and y intertcepts. 

OpenStudy (anonymous):

For inequality 2x + y < 1, the condition will be true for ALL points that are below that line (left side of that line).

OpenStudy (anonymous):